# Nandan Kumar Jha

### v1

@@ 2 Preliminaries
An overview of transformer-based decoder-only architecture. A transformer-based LLM is constructed by sequentially stacking $L$ transformer blocks, where each block is composed of two sub-blocks: an attention mechanism and a feed-forward network (FFN), both having their own residual connections and normalization layers, positioned in the Pre-LN order to improves training stability [25]. Formally, transformer blocks take an input sequence $\mathbf{X}_{\text{in}}\in\mathbb{R}^{T\times d}$, consisting of $T$ tokens of dimension $d$, and transform it into $\mathbf{X}_{\text{out}}$ as follows:
@@ 3 Activation Functions and Their Impact Through Shannon’s Entropy
Well-behaved entropy distribution
Figure 2: Entropy distribution across various configurations of nonlinearities in GPT-2.
We begin by analyzing the headwise entropy distribution of baseline architecture with GELU and ReLU in the FFN, i.e., configurations ${\tt SM+LN+G}$ and ${\tt SM+LN+R}$ respectively. We find that the majority of heads ($\approx$90%) possess entropy values between $\frac{\text{max}}{4}$ and $\frac{\text{3max}}{4}$, where ${\tt max}$ is maximum observed entropy value among all heads (Figure 2). This concentration in the middle entropy range, while avoiding extremes, demonstrates a well-behaved distribution, providing a benchmark for assessing the impact of activation functions on model behavior.
@@
[…]
figureEvaluation loss for baseline and normalization-free GPT-2 models with GELU and ReLU activations.
Configurations Eval PPL +$\Delta$(%)
SM + LN + G 2.69 0.00
SM + LN + R 2.76 2.53
SM + G 3.20 18.92
SM + R 2.94 9.20
[…]
tablePerplexity comparison of baseline and normalization-free GPT-2 models with GELU and ReLU activations in the FFN, trained from scratch. While the baseline model with GELU outperforms its ReLU counterpart, the normalization-free design shows the opposite trend.
Observation 1: ReLU significantly outperforms GELU in LayerNorm-Free LLMs. While GELU is typically preferred over ReLU in conventional transformer-based models due to its smooth and differentiable properties that improve performance and optimization, our empirical findings indicate the opposite trend for LayerNorm-free models— using ReLU in the FFN exhibit better learning dynamics than their GELU counterpart. This leads to an 8.2% improvement in perplexity for GPT-2 (see Figure 3 and Table 3). A similar trend has been observed on the LN-free Pythia-70M model across various context lengths (see Table 2).
@@
To further strengthen our findings, we conducted experiments with a learnable negative slope in the leaky ReLU activation function with two configurations: 1) layer-wise, where each layer has its independent learnable slope, and 2) global, where a single learnable slope is shared across all layers. Results are shown in Figure 3. Interestingly, in the layerwise setting, the early layers initially learn a positive slope while the deeper layers learn a negative slope. However, as training progresses, all layers converge to a near-zero slope. In the global setting, the slope first shifts to positive before converging to near zero. The layerwise entropy dynamics for both cases are plotted in Figure 5.
This highlights the distinct learning dynamics of nonlinearity choices, and a natural preference for zero negative slope, similar to ReLU, in the FFN activation function of the LN-free model.
@@
Figure 4: Evolution of Layerwise entropy when GPT-2 ($L$=12, $H$=12, $d$=768) models with various nonlinearity configurations are trained from scratch on CodeParrot dataset.
@@ 4 Conclusion
In this paper, we investigated the design of LayerNorm-free decoder-only language models and highlighted the critical role of activation functions in such architectures. Our empirical studies revealed that, contrary to conventional practices, the ReLU activation significantly outperforms, an 8.2% improvement in perplexity, the GELU in normalization-free models. We found that models with learnable negative slopes in leaky ReLU activations naturally converge toward zero negative slopes, effectively resembling ReLU. Additionally, we discovered that LayerNorm-free models with GELU activation suffer from entropic overload in early layers, leading to under-utilization of their representational capacity. These findings underscore the necessity of rethinking activation function choices when LayerNorm is absent and suggest that selecting appropriate activations like ReLU enables the development of transformer models that are more efficient, interpretable, and better suited for applications such as private inference and quantization.
@@
Appendix A Why Use Entropy to Evaluate the Impact of Nonlinearities?
@@
Appendix B Design of Experiments
Datasets We train models from scratch using the CodeParrot [24] dataset. The CodeParrot dataset, sourced from 20 million Python files on GitHub, contains 8 GB of files with 16.7 million examples, each with 128 tokens, totaling 2.1B training tokens. We use a tokenizer with a vocabulary of 50K.
Training settings For training on the CodeParrot dataset, we adopt the training settings from [22], which remain consistent across all architectural variations to accurately reflect the impact of the architectural changes. Following [22, 38, 39], all the models are trained on a single RTX 3090 GPU.
B.1 Perplexity as a Reliable Metric to Evaluate the LLMs’ Performance
@@
B.2 Why Training from Scratch to Study Nonlinearities?
Appendix C Additional Results
C.1 Layerwise Entropy Dynamics for Learnable Negative Slope
(a) Layerwise learnable slope
(b) Global slope
Figure 5: Evolution of layerwise entropy in LayerNorm-free GPT-2 models ($L$ = 12, $H$ = 12, $d$ = 768) with a learnable negative slope in the leaky ReLU activation function, trained from scratch on the CodeParrot dataset
C.2 Results on Pythia-70M
Table 2: Results for nonlinearity reduction in Pythia-70M ($L$=6, $H$=8, $d$=512) on CodeParrot dataset. LN-free architecture with ReLU activations in FFN (i.e., ${\tt SM+R}$) significantly outperform their GELU counterpart (i.e., ${\tt SM+G}$) across various content length ($T$).
$T$=128 $T$=256
Eval PPL +$\Delta$(%) Eval PPL +$\Delta$(%)
SM+LN+G 3.512 0.00 3.054 0.00
SM+LN+R 3.590 2.22 3.107 1.73
SM+G 4.086 16.35 3.570 16.87
SM+R 3.736 6.36 3.273 7.17
@@ Appendix D Future Work
Longitudinal studies and downstream task performance To validate the long-term viability of this approach, conducting longitudinal studies is crucial. Ensuring that the performance benefits observed with ReLU activation persist over extended training periods and across different tasks would reinforce our findings. Furthermore, while our study focused on perplexity and entropy metrics, future work should analyze how the choice of activation function affects performance on various downstream tasks. Investigating the implications for fine-tuning processes could also provide valuable insights for practical applications.
Hybrid activation strategies Exploring hybrid activation strategies presents another promising avenue. By using different activation functions in different parts of the model—such as combining ReLU in the earlier layers with GELU in the later layers—we could strike a balance between the benefits observed in our study and the traditional advantages of GELU. This approach may enhance model performance while maintaining computational efficiency.
Interpretability and practical applications Given the observed differences in entropy distribution between ReLU and GELU models, future research could explore how different activation functions impact the interpretability of LayerNorm-free models. This could potentially lead to more explainable AI systems, addressing a critical need in the field. Integrating these findings into practical applications like private inference and quantization is also promising, as it could improve both model efficiency and security.

### v2

@@ 2 Preliminaries
2.1 Overview of Transformer-based Decoder-only Architectures
A transformer-based LLM is constructed by sequentially stacking $L$ transformer blocks, where each block is composed of two sub-blocks: an attention mechanism and a feed-forward network (FFN), both having their own residual connections and normalization layers, positioned in the Pre-LN order to improves training stability [25]. Formally, transformer blocks take an input sequence $\mathbf{X}_{\text{in}}\in\mathbb{R}^{T\times d}$, consisting of $T$ tokens of dimension $d$, and transform it into $\mathbf{X}_{\text{out}}$ as follows:
@@ 2.2 Entropy as a Metric for Attention Score Distribution
2.3 Dataset and Training Methodology
We train our models from scratch using the CodeParrot dataset [24], a standard benchmark for LLMs [22, 16]. The dataset, derived from 20 million Python files on GitHub, consists of 8 GB of data with 16.7 million examples, each containing 128 tokens, amounting to a total of 2.1 billion training tokens. For tokenization, we utilize a tokenizer with a vocabulary size of 50K.
For training on the CodeParrot dataset, we adopt the settings from [22], ensuring consistency across all architectural variations to isolate the effects of the changes. In line with prior works [22, 34, 35], all models are trained using a single RTX 3090 GPU.
@@ 3 Activation Functions and Their Impact Through Shannon’s Entropy
Well-behaved entropy distribution We begin by analyzing the headwise entropy distribution of baseline architecture with GELU and ReLU in the FFN, i.e., configurations ${\tt SM+LN+G}$ and ${\tt SM+LN+R}$ respectively. We find that the majority of heads ($\approx$90%) possess entropy values between $\frac{\text{max}}{4}$ and $\frac{\text{3max}}{4}$, where ${\tt max}$ is maximum observed entropy value among all heads (Figure 2a). This concentration in the mid-entropy range, avoiding extremes, demonstrates a well-behaved distribution, providing as a benchmark for assessing the impact of architectural modifications, such as activation function simplification, on model behavior.
@@
(a) Headwise entropy distribution
(b) Evaluation loss curve
Figure 2: Headwise entropy distribution and evaluation loss for baseline and normalization-free GPT-2 models, using GELU and ReLU activations, trained from scratch on CodeParrot dataset.
Observation 1: ReLU significantly outperforms GELU in LayerNorm-Free LLMs. While GELU is typically preferred over ReLU in conventional transformer-based models due to its smooth and differentiable properties that improve performance and optimization, our empirical findings indicate the opposite trend for LayerNorm-free models— using ReLU in the FFN exhibit better learning dynamics than their GELU counterpart. This leads to an 8.2% improvement in perplexity for GPT-2 (see Figure 2 and Table 2). A similar trend has been observed on the normalization-free Pythia-70M model across various context lengths.
Table 2: Perplexity comparison between baseline and normalization-free GPT-2 ($L$=12, $H$=12, $d$=768) and Pythia-70M ($L$=6, $H$=8, $d$=512) models, using GELU and ReLU activations in the FFN, trained from scratch on CodeParrot dataset. While GELU outperforms ReLU in baseline models, the normalization-free models exhibit the opposite trend.
GPT-2 ($T$=128) Pythia-70M ($T$=128) Pythia-70M ($T$=256)
Eval PPL +$\Delta$(%) Eval PPL +$\Delta$(%) Eval PPL +$\Delta$(%)
SM+LN+G 2.688 0.00 3.512 0.00 3.054 0.00
SM+LN+R 2.757 2.53 3.590 2.22 3.107 1.73
SM+G 3.197 18.92 4.086 16.35 3.570 16.87
SM+R 2.936 9.20 3.736 6.36 3.273 7.17
To further strengthen our findings, we conducted experiments with a learnable negative slope in the leaky ReLU activation function with two configurations: 1) layer-wise, where each layer has its independent learnable slope, and 2) global, where a single learnable slope is shared across all layers. Interestingly, in the layerwise setting, the early layers initially learn a positive slope while the deeper layers learn a negative slope. However, as training progresses, all layers converge to a near-zero slope. In the global setting, the slope first shifts to positive before converging to near zero (see Figure 3).
@@
When comparing the layerwise entropy dynamics in both cases (Figure 4e and Figure 4f) with the normalization-free model using ReLU activations (Figure 4d), we observed near-identical patterns. This highlights the a natural preference for zero negative slope, similar to ReLU, in the FFN activation function of the normalization-free model.
@@
(e) SM + LearnableNegativeSlope (Layerwise)
(f) SM + LearnableNegativeSlope (Global)
Figure 4: Evolution of Layerwise entropy when GPT-2 ($L$=12, $H$=12, $d$=768) models with various nonlinearity configurations are trained from scratch on CodeParrot dataset. Evolution of layer-wise entropy during training of GPT-2 models ($L$=12, $H$=12, $d$=768) with different nonlinearity configurations on the CodeParrot dataset. The near-identical entropy dynamics in Figures d, e, and f underscore a natural preference for a zero negative slope, similar to ReLU, in the FFN activation function of the normalization-free model.
@@
(a) NaNs observed with a negative slope of $1e$-2
(b) Entropy dynamics with a negative slope of $1e$-2
(c) NaNs observed with a negative slope of $5e$-2
(d) Entropy dynamics with a negative slope of $5e$-2
(e) NaNs observed with a negative slope of $1e$-1
(f) Entropy dynamics with a negative slope of $1e$-1
(g) NaNs observed with a negative slope of $2e$-1
(h) Entropy dynamics with a negative slope of $2e$-1
Figure 5: Training instability, indicated by NaNs, and corresponding entropy dynamics in Normalization-Free GPT-2 ($L$=12, $H$=12, $d$=768) models with fixed negative slopes in the leaky ReLU. The larger the negative slope, the earlier the training instability and entropy collapse occurred.
Training instability and entropy dynamics with a fixed negative slope In normalization-free LLMs, ReLU-like activation functions, with a near-zero negative slope, naturally stand out as a preferred choice compared to the conventional GELU, offering both improved predictive performance and stable training dynamics. This makes exploring fixed negative slopes in leaky ReLU activations particularly intriguing.
To systematically investigate this, we conducted a series of experiments on normalization-free GPT-2 models, adjusting the negative slopes to fixed values of $1e$-2, $5e$-2, $1e$-1, and $2e$-1. We assessed training instability by monitoring the frequency and distribution of NaN values across model layers and evaluated entropy dynamics across the model’s depth. The results are shown in Figure 5.
For a negative slope of $1e$-2, we observed sporadic occurrences of NaNs primarily in the last layer (Layer 11), as shown in Figure 5a, with no entropy collapse (Figure 5b). However, as the negative slope increased to $5e$-2, $1e$-1, and $2e$-1, NaNs began to appear consistently in deeper layers, as evidenced by the NaN counts in Figures 5c, 5e, and 5g, respectively. These consistent NaNs correlated with entropy collapses in the deeper layers, indicating a strong relationship between increased negative slope and training instability (Figures 5d, 5f, and 5h).
An interesting trend emerged: the larger the negative slope, the earlier the training instability and entropy collapse occurred. For example, with a slope of $2e$-1, instability occurred almost immediately (Figure 5g), and entropy collapses in deeper layers (Figure 5h) occurred much sooner compared to lower negative slopes. This suggests that as the slope increases, the window of stable training narrows, making it crucial to choose the appropriate negative slope to avoid early instability and entropy collapse in normalization-free models.
Broader implications of activation function characteristics in normalization-free models The emergence of training instability and entropy collapse at larger negative slopes underscores the sensitivity of normalization-free LLMs to the choice of activation function parameters. The strong correlation between larger negative slopes and earlier training instability suggests that even seemingly minor changes to the negative slope of the leaky ReLU function can significantly influence the model’s ability to maintain stability during training. Specifically, larger negative slopes in leaky ReLU activations aggravate the proliferation of NaNs and exacerbate entropy collapse in deeper layers, leading to earlier and more pronounced training instability.
This suggests that a near-zero negative slope strikes a crucial balance in normalization-free LLMs, offering sufficient nonlinearity while maintaining training stability.
@@ 4 Conclusion
In this paper, we investigated the design of normalization-free decoder-only language models and highlighted the critical role of activation functions in such architectures. Our empirical studies revealed that, contrary to conventional practices, the ReLU activation significantly outperforms, an 8.2% improvement in perplexity, the GELU in normalization-free models. We found that models with learnable negative slopes in leaky ReLU activations naturally converge toward zero negative slopes, effectively resembling ReLU.
Additionally, we discovered that LayerNorm-free models with GELU activation suffer from entropic overload in early layers, leading to under-utilization of their representational capacity. These findings underscore the necessity of rethinking activation function choices when LayerNorm is absent and suggest that selecting appropriate activations like ReLU enables the development of transformer models that are more efficient, interpretable, and better suited for applications such as private inference and quantization.
@@
Notes. This workshop submission delves into one of the key findings—the LayerNorm-free design–from our comprehensive paper AERO: Softmax-Only LLMs for Efficient Private Inference. The code and implementation are available at https://github.com/Nandan91/relu-revival-normfree.
@@
Appendix A Why Training from Scratch to Study Nonlinearities?
Appendix B Why Use Entropy to Evaluate the Impact of Nonlinearities?
@@
Appendix C Perplexity as a Reliable Metric to Evaluate the LLMs’ Performance
@@ Appendix D Future Work
Downstream task performance While our study focused on perplexity and entropy metrics, future work should analyze how the choice of activation function affects performance on various downstream tasks. Investigating the implications for fine-tuning processes could also provide valuable insights for practical applications.
Hybrid activation strategies Exploring hybrid activation strategies presents another promising research direction. By using different activation functions in different parts of the model—such as combining GELU in the earlier layers with ReLU in the later layers—we could strike a balance between the benefits observed in our study and the traditional advantages of GELU. This approach may enhance model performance while maintaining computational efficiency.
Interpretability and practical applications Given the observed differences in entropy distribution between ReLU and GELU models, future research could explore how different activation functions impact the interpretability of LayerNorm-free models. This could potentially lead to more explainable model design, addressing a critical need in the field. Integrating these findings into practical applications like private inference and quantization is also promising, as it could improve both model efficiency and security.